OpenStudy (anonymous):

one fifth of the movie audience left the theater in dusgust. If 66 angry patrons walked out, how many were there originally?

OpenStudy (nebi):

66.2???? i think

OpenStudy (jamesj):

@JV: this is like one of the problems we worked a couple of hours ago.

OpenStudy (jamesj):

The one with 90%. So follow that procedure, or the one across is recommending.

OpenStudy (anonymous):

@ James, Yeah I know but I still don't understand how to do the problem... :/

OpenStudy (across):

Nebi, :) I wonder what 66.2 people look like!

OpenStudy (jamesj):

Let X be the number of people in the theater originally. How do write one fifth of X?

OpenStudy (anonymous):

20(5)=100

OpenStudy (anonymous):

66(5)=?

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 5\end{matrix}\right)x\]

OpenStudy (anonymous):

?

OpenStudy (jamesj):

Do you mean X/5 ?

OpenStudy (anonymous):

I'm getting extremely confused

OpenStudy (jamesj):

X/5 is one fifth of X. Do agree with that?

OpenStudy (anonymous):

If 66 people = 1/5 (one fifth).....then 66 x 5 = 5/5 66 x 5 = 330

OpenStudy (anonymous):

sure I guess... Idk what it means but I'm sure your right James

OpenStudy (anonymous):

i just typed that

OpenStudy (jamesj):

Well, it's true that X/5 means one fifth of X Now what do we know about one fifth of X, the number of patrons? It is equal to 66 That is: X/5 = 66 How do you now solve that equation for X?

OpenStudy (anonymous):

I found that it was 330... Sorry to make you upset?......

OpenStudy (anonymous):

Why did you delet your comment?

OpenStudy (jamesj):

Yes, 330 is right.

OpenStudy (anonymous):

Or you can do it in terms of percentages: 66 is 20%. We know this because it tells us that 66 is one fifth of the audience. So, you do 66 divided by 0.20 (the 20%) = 330 You can check its right by reversing it....330 x 0.2 = 66 !